# Misinformation is not about Bad Facts: An Analysis of the Production and Consumption of Fringe Content

### v1

@@ Abstract
What if misinformation is not an information problem at all? Our findings suggest that online fringe ideologies spread through the use of content that is consensus-based and “factually correct”. We found that Australian news publishers with both moderate and far-right political leanings contain comparable levels of information completeness and quality; and furthermore, that far-right Twitter users often share from moderate sources. However, a stark difference emerges when we consider two additional factors: 1) the narrow topic selection of articles by far-right users, suggesting that they cherrypick only news articles that engage with specific topics of their concern, and 2) the difference between moderate and far-right publishers when we examine the writing style of their articles. Furthermore, we can even identify users prone to sharing misinformation based on their communication style. These findings have important implications for countering online misinformation, as they highlight the powerful role that users’ personal bias towards specific topics, and publishers’ writing styles, have in amplifying fringe ideologies online.
@@ 1 Introduction
Misinformation can broadly be defined as false and inaccurate narratives that people may believe and communicate despite the existence of a social or expert consensus on the topic. Although not purposefully spread to create harm (which would be called disinformation), misinformation can cause significant societal disruption. For example, anti-vaccination rhetoric can cause families to withhold vaccination and expose children to life-threatening but preventable diseases such as measles and poliomyelitis, with potentially fatal consequences (Attwell, Smith, and Ward 2021). The reach and consequences of misinformation are amplified by the increasing use of online social media as a primary source of news and information (Shearer and Mitchell 2021). For example, Kydd (2021) blames social media for mainstreaming conspiracy theorizing, resulting in the attack on the US Capitol by devout QAnon followers. Due to the diverse range of ways misinformation can influence believers, it follows that misinformation cannot be studied in isolation. It is crucial to consider what is being shared in context: who is sharing it, who are the likely consumers and how it is being shared to reach the audiences effectively. In this work, we examine how far-right and fringe online groups share and leverage news media articles to advance their narratives, to understand the role of news publishers in potentially unintentionally propagating misinformation.
Figure 1: Summary of the methods: we start by analysing contents of Australian news articles and sharing patterns of those articles in Twitter environment. Then we use linguistic styles captured by LIWC, Grievance dictionary and StyloMetrix, instead of contents, to identify extreme groups in Facebook. Next, we identify text styles employed by the extreme groups and classify them. Finally, we distinguish misinformation production and consumption using styles.
Research Questions: In this paper, we use content- and style-based measures to show that misinformation, like any other content, can be styled to target particular online cohorts. Together, these elements and their users make up a misinformation ecosystem. In particular, we answer the following research questions (RQs):
  • RQ1: What role do the style and content of news media play in enabling the spread of misinformation?
@@
Our contributions are as follows:
@@
    We propose a classifier to identify linguistic styles within extreme online groups as well as styles that are commonly used in the online misinformation space. (RQ2).
@@
    We show that production and consumption of misinformation exhibit different patterns that can be differentiated by styles (RQ3).
Fig. 1 depicts the steps of our analysis, the used datasets and methods. We expand each contribution below.
Landscape of Australian news (Section 3)
Starting from the assumption that misinformation is an information production problem, much of the prior work has studied the connection between misinformation, and the publishers’ political leaning (Calvillo et al. 2020; Das and Ahmed 2022). We conducted a comprehensive analysis of Australian news media consumption, focusing on several key aspects. Firstly, we utilized the Trust Index, a metric measuring information completeness, to compare news articles from legacy news media publishers with distinct ideological stances. Secondly, we examined writing style differences across these sources. Additionally, we delved into the sharing patterns of news articles by highly partisan online users.
Our analysis revealed content-based differentiations among news sources and extreme ideological cohorts. Interestingly, we observed significant differences when comparing articles shared by extreme online users with articles published by far-right news media, particularly in terms of linguistic patterns. Notably, we found less distinction between far-right and moderate articles. Even in cases where differences were statistically significant, the effect sizes were small, indicating that the sharing of extreme content is less tied to the source’s reliability and more about how users interpret and leverage a source.
This leads us to the conclusion that misinformation may be less of an issue related to information production and more of an issue tied to information consumption. To explore this further, we investigate whether it is an information consumption problem.
Classification of Extreme Online Groups and Styles (Section 4)
This study uses stylistic metrics from established dictionaries (such as LIWC and the Grievance dictionary) to construct stylistic classifiers. Using these stylistic classifiers, we can accurately distinguish vulnerable misinformation consumers (like anti-vaccination and far-right) from regular online social media users (Section 4.1). We further show that we can detect extreme ideological users’ writing styles, as categorized by our manual labeling (Section 4.2). Lastly, we show that we can successfully distinguish far-right sources from far-right consumers using a stylistic classifier (Section 4.3).
@@ 2.1 News and social media datasets
To evaluate the potential role of Australian news publications in facilitating misinformation dissemination, we collect Australian news articles sourced from Google News via The Daily Edit (TDE) platform. TDE aggregates news articles from Google News, encompassing 14 distinct topics: ‘Climate Change,’ ‘Sport,’ ‘Human Migration,’ ‘World,’ ‘Finance,’ ‘Technology,’ ‘Taiwan,’ ‘Top Stories,’ ‘Entertainment,’ ‘Australia,’ ‘Business,’ ‘Health,’ ‘Science,’ and ‘China.’ These are all the topics available when the region is set to Australia (AU) in Google News. The collection period for the news articles spans from November 9, 2022, to January 11, 2023. The dataset contains $214,754$ articles published by $9,929$ news publishers.
TDE computes a Trust Index score for each article as follows. Articles about the same event are grouped into stories. Articles are represented as sequences of sentences; all similar sentences within a story are clustered together, and clusters are interpreted as details. Consequently, each article has a set of supporting details (common narrative elements across multiple sources in a story). TDE computes an article’s Trust Index based on the percentage of details it covers from a story – i.e., the article’s informational completeness.
@@
We use an external media bias dataset from allsides [^1] https://www.allsides.com/, which assesses the political leaning of $473$ news publishers on a five-point scale, ranging from extreme- and moderate-left to center to moderate- and extreme-right. Note that these media bias ratings represent the average viewpoint of individuals across the political spectrum rather than the perspective of any single individual or group (Park et al. 2021; Newman et al. 2021). We consolidate the extreme- and moderate-left categories into a single ‘left’ class and merge the extreme- and moderate-right categories into a ‘right’ class. This results in three political leaning classes: left, center, and right.
@@
We use a manually curated list containing $1,496$ Australian far-right coded Twitter users (Bailo, Johns, and Rizoiu 2023) and collect their $3,665,809$ most recent tweets using the Twitter API (at most $3,200$ tweets for each user). The collected tweets span from February 22, 2009 to December 9, 2022. These tweets contained $1,827,162$ urls and among the urls, $11,643$ articles were also in Google News dataset (Section 3). We expand the articles by linking to the publishers present in either the Google News or allsides dataset which results in $215,242$ news articles (Section 4). Then we collected the content of each article using urllib.parse Python library222 https://docs.python.org/3/library/urllib.parse.html. We successfully downloaded the texts of $155,669$ articles at the end.
@@
We manually assemble two lists of Facebook pages for specific ideologies, namely Australian Antivax and Far-right pages. Australian Far-right has 14 Facebook pages and Australian Antivax has 6 Facebook pages. Using CrowdTangle API, we collect all posts from the Facebook pages between January 23, 2021 and February 24, 2023. We collect $6017$ posts from the Far right pages and $2969$ posts from the Antivax pages.
@@ 2.2 Linguistic Measurements of Style
is one of the most widely used text analysis tools in psychology, which has recently been adopted by computational social scientists to draw insights into human behavior through computational methods (Chung and Pennebaker 2018). It includes both content (e.g., death, religion) and function (e.g., conjunctions, articles) words. LIWC (version 2022) [^3] https://www.liwc.app/ has 117 categories in total. However, in order to capture the stylistic differences of the extreme groups without contamination from content, we removed content-related features such as religion and family. We use 89 style-related categories in Section 4.
@@
is a psycholinguistic dictionary that can be used to automatically understand language use in the context of grievance-fueled violence threat assessment (van der Vegt et al. 2021). We use it because a study about militant right-wing extremism (Stankov 2021) suggests that a critical ingredient of a militant extremist mindset distinguishing it from social conservatism is grudge, which it can detect. We include Grievance dictionary even though it is a content-based measurement due to its ability to extract violence and threat-specific words.
@@
is a grammar-related statistical representation of text. It allows for representing a text sample of any length with a linguistic vector of a fixed size (Okulska and Zawadzka 2023). It has several preferable characteristics over well-known contextual embedding, such as BERT. First, StyloMetrix vectors encode entire documents, resolving the issue of varying text lengths. This could help when combining texts from multiple platforms, such as Facebook and Twitter. Second, StyloMetrix vectors encode the entire sample’s stylistic structure, not the words’ “meanings”. Okulska and Zawadzka (2023) show that show that StyloMetrix (without a semantic layer) is sufficient to detect the genre of the text.
3 Australian News: production and sharing (RQ1)
Here, we answer RQ1 and investigate whether the content from publishers with different ideological leanings exhibits deviation in terms of topic coverage, the Trust Index of articles published, and linguistic patterns. We also examine how users share these articles to promote their own agendas.
@@ 3.1 Coverage and Trust of publishers
We extracted the $41,022$ Google News articles from publishers with identified stances based on the media bias data (note that not all publishers from the Google News dataset are present in the allsides dataset). Table 1 shows the number of articles from the stance-identified publishers for each topic. Across all topics, except ‘Entertainment,’ we observe a higher number of left-leaning articles than right-leaning ones. Particularly noteworthy is the disparity in percentages within the ‘Climate change’ and ‘Technology’ topics, where the count of right-leaning articles is substantially lower than that of left-leaning articles.
@@
We conduct a comparative analysis of the Trust Index for articles from left-, center-, and right-leaning publishers. For each topic and political group, Table 2 provides the mean ($\mu$), standard deviation ($\sigma$), and the number of articles ($N$). We assess whether there is a significant difference between Trust Index of left- and right-leaning news articles using independent samples $t$-tests. We report the test p-value, and we emphasize statistically significant results ($p<0.05$). We also quantify the effect size for each test using Cohen’s d (Cohen 2013). Generally, a d value of 0.2 indicates a small effect size and $0.5$ is considered medium.
There are several observations from Table 2. First, the mean Trust Index for articles from left-leaning publishers (L in Table 2) consistently surpasses that of the right-leaning, regardless of the topic’s statistical significance. This indicates that, in general, left-leaning articles tend to be more informationally complete. The standard deviations of the Trust Index is nearly identical across all groups and topics. Second, several topics such as ‘World’, ‘China’, ‘Technology’, ‘Finance’ and ‘Human Migration’ achieve statistical significance but only moderate effect size. ‘Taiwan’ exhibits the largest effect size among all topics; however, it does not reach statistical significance due to the limited number of articles in the ’Taiwan’ category (refer to Table 1). Conversely, ‘Australia’ shows a statistically significant difference, albeit with only a small effect size ($0.11$).
@@ Conclusion.
While the distinction between left- and right-leaning publishers is statistically significant for specific topics, the practical impact of this difference is marginal in the real world, given the small or non-existent effect size.
@@ 3.2 Is the content from far-right publishers different from the moderates?
Here, we ask whether Google News articles’ content and writing style differ based on the political leaning of the publishers. More specifically, we are interested in far-right publishers (as per the allsides classification) vs all the others that we dub here moderates.
@@
We manually inspect randomly selected articles and identify four specific topics – ‘Top Stories,’ ‘Australia,’ ‘Finance,’ and ‘Climate Change’ – that exhibited interesting content contrasts. Table 3 shows the number of articles from far-right and moderate publishers for each category. We gather the content of each article using the beautifulsoup Python library444beautifulsoup: https://pypi.org/project/beautifulsoup4/. Next, we use LIWC to analyze the text of each article and perform t-tests to find whether articles from far-right publishers show statistically significant differences from moderates. To present significant results, we use the Bonferroni correction, i.e., $p<\cfrac{\alpha}{117}$ where $\alpha=0.05$.

### v2

@@ Abstract
What if misinformation is not an information problem at all? To understand the role of news publishers in potentially unintentionally propagating misinformation, we examine how far-right and fringe online groups share and leverage established legacy news media articles to advance their narratives. Our findings suggest that online fringe ideologies spread through the use of content that is consensus-based and “factually correct”. We found that Australian news publishers with both moderate and far-right political leanings contain comparable levels of information completeness and quality; and furthermore, that far-right Twitter users often share from moderate sources. However, a stark difference emerges when we consider two additional factors: 1) the narrow topic selection of articles by far-right users, suggesting that they cherry pick only news articles that engage with their preexisting worldviews and specific topics of concern, and 2) the difference between moderate and far-right publishers when we examine the writing style of their articles. Furthermore, we can identify users prone to sharing misinformation based on their communication style. These findings have important implications for countering online misinformation, as they highlight the powerful role that personal biases towards specific topics and publishers’ writing styles have in amplifying fringe ideologies online.
@@ 1 Introduction
Misinformation has historically been understood as an information or factual accuracy issue, where concerns arise when inaccurate narratives emerge despite the existence of a social or expert consensus on the topic. This has led to a focus on interventions that seek to remedy inaccuracies, like fact-checking, presuming that people will willingly change their perspective on an issue when presented with an alternative, authoritative form of information. These approaches, however, overlook the human side of the problem, such as the everyday anxieties and grievances that can motivate belief in misinformation (Booth et al. 2024). The ubiquity of digital technologies and social media means the reach and consequences of misinformation are increasingly amplified (Attwell, Smith, and Ward 2021; Shearer and Mitchell 2021; Kydd 2021).
Technological interventions by social media platforms have thus far proved to be limited in controlling the spread of misinformation (Johns et al. 2024). As a result, it follows that we must turn our attention to the heart of the issue: that humans have uncertainties and worries, and seek solace through answers. Thus, when considering misinformation online, it is crucial to consider what is being shared not as an abstract issue of factual inaccuracy, but in its context: who is sharing it, who are the likely consumers, and how is it being shared to reach audiences effectively.
To understand the role of news publishers in potentially unintentionally propagating misinformation, we examine how far-right and fringe online groups share and leverage established legacy news media articles to advance their narratives. We also demonstrate that far-right and moderate news articles in our sample differ not in the information completeness of the articles, but in their writing style, thus providing a new perspective on misinformation as an issue of style. From this, we consider how style might be leveraged by moderate and consensus-based news sources to better counter the influence of online misinformation.
Research Questions. We use content- and style-based measures to show that misinformation, like any other content, can be styled to target particular online cohorts. In particular, we answer the following research questions (RQs):
  • RQ1: What role does the style and content of news media play in enabling the spread of misinformation?
@@
Figure 1: Summary of the methods: we start by analysing contents of Australian news articles and sharing patterns of those articles in Twitter environment. Then we use linguistic styles – not the actual article contents – measured with LIWC, Grievance dictionary and StyloMetrix to identify extreme groups in Facebook. Next, we identify text styles employed by the extreme groups and classify them. Finally, we distinguish misinformation production and consumption using styles.
The graphic in Figure 1 depicts the steps of our analysis, the datasets used and the methods. We begin by analyzing the content completeness of news articles based on their political ideology. We compare 1) publishers of different ideologies and 2) production and consumption by far-right users. Based on the findings that content completeness of news articles does not significantly differ based on publishers’ ideology, we leverage the text style (instead of its content) to investigate differences and distinguish between publishers of different ideologies and fringe online communities. Lastly, we show that far-right production and consumption can be distinguished based on the employed writing style. We expand on the details of our contributions below.
We conducted a comprehensive analysis of Australian news media consumption, focusing on several key aspects. Firstly, we utilized the Trust Index (The Daily Edit, Inc 2023) – a metric measuring information completeness – to compare news articles from legacy news media publishers to those with distinct ideological stances. Secondly, we examined the differences in writing style across these sources. Additionally, we delved into the sharing patterns of news articles by highly partisan online users. This differs from much of the prior work that has started from the assumption that misinformation is an information problem, and studied the connection between misinformation and the publishers’ political leaning (Calvillo et al. 2020; Das and Ahmed 2022).
Our analysis revealed a distinct difference between what articles were published by far-right news sources and what articles were shared by the extreme (here, also far-right) ideological cohort in our sample, particularly in terms of linguistic patterns. However, we notably found less distinction between far-right and moderate articles. Even in cases where differences were statistically significant, the effect sizes were small, indicating that the sharing of extreme content is less tied to the source’s perceived ideological stance and more about how users interpret the value of an article as a tool to advance their own worldview. This is to say that users will individually cherry pick articles to share on the basis of their own views, regardless of what source published it.
This leads us to the conclusion that misinformation is not strictly a matter of factually-inaccurate content being spread by far-right sources, but that mostly-accurate content can be selectively consumed to fuel an existing fringe worldview. To explore this further, we investigate how such groups express their worldviews through different linguistic styles.
This study uses stylistic metrics from established dictionaries (such as LIWC and the Grievance dictionary) to construct stylistic classifiers. Using these stylistic classifiers, we can accurately distinguish consumers who may be vulnerable to misinformation (like anti-vaccination and far-right sentiment) from regular online social media users. We further show that we can detect extreme ideological users’ writing styles, as categorized by our manual labeling. Lastly, we show that we can successfully distinguish far-right sources from far-right consumers using a stylistic classifier.
The summary of our contributions are as follows:
@@
    We propose a classifier to identify linguistic styles within extreme online groups as well as styles that are commonly used in the online misinformation space (RQ2).
@@
    We show that the production and consumption of misinformation exhibit different patterns that can be differentiated by styles (RQ3).
@@ 2.1 News and social media datasets
To evaluate the potential role of Australian news publications in facilitating misinformation dissemination, we collected Australian news articles sourced from Google News via The Daily Edit (TDE) platform. TDE aggregates news articles from Google News, encompassing 14 distinct topics: ‘Climate Change,’ ‘Sport,’ ‘Human Migration,’ ‘World,’ ‘Finance,’ ‘Technology,’ ‘Taiwan,’ ‘Top Stories,’ ‘Entertainment,’ ‘Australia,’ ‘Business,’ ‘Health,’ ‘Science,’ and ‘China.’ This is the exhaustive list of topics available when the region is set to Australia (AU) in Google News. The published period for the news articles spans from March 1, 2014, to January 12, 2023. The dataset contains $214,754$ articles published by $9,929$ news publishers.
TDE computes a Trust Index score for each article as follows. Articles about the same event are grouped into stories. Articles are represented as sequences of sentences; all similar sentences within a story are clustered together, and clusters are interpreted as details. Consequently, each article has a set of supporting details (common narrative elements across multiple sources in a story). TDE computes an article’s Trust Index based on the percentage of details it covers from a story – the article’s informational completeness.
@@
We use an external media bias dataset from allsides [^1] https://www.allsides.com, which assesses the political leaning of $473$ news publishers on a five-point scale, ranging from extreme- and moderate-left to center to moderate- and extreme-right. These media bias ratings represent the average viewpoint of individuals across the political spectrum rather than the perspective of any single individual or group (Park et al. 2021; Newman et al. 2021). We consolidate the extreme- and moderate-left categories into a single ‘left’ class and merge the extreme- and moderate-right categories into a ‘right’ class. This results in three political leaning classes: left, center, and right. Later, we also examine articles from publishers belonging to the ‘extreme-right’ range by allsides.
@@
We use a manually curated list containing $1,496$ Australian far-right coded Twitter users (Bailo, Johns, and Rizoiu 2023) and collect their $3,665,809$ most recent tweets using the Twitter API (at most $3,200$ tweets for each user). The collected tweets span from February 22, 2009 to December 9, 2022. These tweets contained $1,827,162$ urls consisting of $11,643$ articles in Google News dataset (Section 3). We expand the articles by linking to the publishers present in either the Google News or allsides dataset which results in $215,242$ news articles (Section 4). We scraped the content of each article using urllib.parse Python library222 https://docs.python.org/3/library/urllib.parse.html. We successfully downloaded the texts of $155,669$ articles at the end because the rest of the links were not available.
@@
We manually assemble two lists of Facebook pages for specific ideologies, namely Australian Antivax and Far-right pages. Australian Far-right has $14$ Facebook pages and Australian Antivax has six Facebook pages. Using the CrowdTangle API, we collect all posts from the Facebook pages between January 23, 2021 and February 24, 2023. We collect $6,017$ posts from the Far right pages and $2,969$ posts from the Antivax pages.
@@ 2.2 Linguistic Measurements of Style
is one of the most widely used text analysis tools in psychology, which has recently been adopted by computational social scientists to draw insights into human behavior through computational methods (Chung and Pennebaker 2018). This tool captures words relating to content (e.g., death, religion) and function (e.g., conjunctions, articles). LIWC (version 2022) [^3] https://www.liwc.app/ has $117$ categories in total. However, to capture the extreme groups’ stylistic differences without contamination from content, we removed content-related features (such as religion, family and others – see Appendix A), yielding a total of $89$ style-related categories.
@@
is a psycholinguistic dictionary to capture language use in the context of grievance-fueled violence threat assessment (van der Vegt et al. 2021). Grudge has been shown as the critical ingredient that distinguishes militant extremist mindset from social conservatism (Stankov 2021); grudge (alongside confusion) is also one of the ingredients of misinformation consumption (Booth et al. 2024). We include Grievance dictionarydue to its ability to extract violence and threat-specific words.
@@
is a grammar-related statistical representation of text. This tool allows for representing a text sample of any length with a linguistic vector of a fixed size (Okulska and Zawadzka 2022) offering several preferable characteristics over well-known contextual embedding, such as BERT. First, StyloMetrix vectors encode entire documents, resolving the issue of varying text lengths. This could help when combining texts from multiple platforms, such as Facebook and Twitter. Second, StyloMetrix vectors aim to encode the entire sample’s stylistic structure, not the words’ meanings.
3 The Content: News Production and Sharing
Here, we answer RQ1 and demonstrate that there is only a marginal difference in the information completeness of left- and right-leaning publishers, with the left-leaning publishers publishing slightly more articles. Furthermore, there are few differences in linguistic patterns between far-right and moderate publishers. However, we see a marked difference when comparing what far-right publishers produce compared to what far-right users actually share online.
@@ 3.1 Coverage and Trust of publishers
We extracted the $40,922$ Google News articles from publishers with identified stances based on the media bias data (note that not all publishers from the Google News dataset are present in the allsides dataset, and we exclude any missing publishers). Table 1 shows the number of articles from the stance-identified publishers for each topic. Across all topics, except ‘Entertainment,’ we observe a higher number of left-leaning articles than right-leaning. Particularly noteworthy is the disparity in percentages within the ‘Climate change’ and ‘Technology’ topics, where the count of right-leaning articles is substantially lower than that of left-leaning articles. This indicates that the Australian news media is generally perceived to be more left-leaning.
@@
We conduct a comparative analysis of the Trust Index for articles from left-, center-, and right-leaning publishers. For each topic and political group, Table 2 provides the mean ($\mu$), standard deviation ($\sigma$), and the number of articles ($N$). We assess whether there is a significant difference between Trust Index of left- and right-leaning news articles using independent samples $t$-tests. We report the test p-value, and we emphasize statistically significant results ($p<0.05$). We also quantify the effect size for each test using Cohen’s d (Cohen 2013). Generally, a d value of $0.2$ indicates a small effect size and a value of $0.5$ is considered a medium effect.
There are several observations from Table 2. First, the mean Trust Index for articles from left-leaning publishers (L in Table 2) consistently surpasses that of the right-leaning, regardless of the topic’s statistical significance. This indicates that, in general, left-leaning articles tend to be more informationally complete. The standard deviations of the Trust Index are nearly identical across all groups and topics. Second, several topics such as ‘World’, ‘China’, ‘Technology’, ‘Finance’ and ‘Human Migration’ achieve statistical significance but only moderate effect size. ‘Taiwan’ exhibits the largest effect size among all topics; however, it does not reach statistical significance due to the limited number of articles in the ‘Taiwan’ category (refer to Table 1). Conversely, ‘Australia’ shows a statistically significant difference, albeit with only a small effect size (Cohen’s $d=0.11$).
@@ Conclusion.
While the distinction between left- and right-leaning publishers’ information completeness is statistically significant for specific topics, the practical impact of this difference is marginal in the real world, given the small or non-existent effect size. This is unexpected, as we might assume fringe news outlets to consistently include misinformation or factual inaccuracies.
@@ 3.2 Is the content from far-right publishers different from the moderates?
Here, we demonstrate there is insufficient evidence using LIWC to show that Google News articles’ content and writing style differ based on the political leaning of the publishers. More specifically, we find this conclusion is consistent with far-right publishers (‘extreme-right’ per allsides classification) compared to all the others that we group here as “moderates”.
@@
We qualitatively inspected randomly selected articles and identified four specific topics – ‘Top Stories,’ ‘Australia,’ ‘Finance,’ and ‘Climate Change’ – that exhibited interesting content contrasts. Table 3 shows the number of articles from far-right and moderate publishers for each category. We gather the content of each article using the beautifulsoup Python library444beautifulsoup: https://pypi.org/project/beautifulsoup4/. Next, we use LIWC to analyze the text of each article and perform t-tests to find whether articles from far-right publishers show statistically significant differences from moderates. To present significant results, we use the Bonferroni correction, i.e., $p<\alpha/117$ where $\alpha=0.05$ (note: there are $117$ categories in LIWC, see Section 2.2).